Exercise 10-2 Straight-Line: Amortization of bond discount LO P2 Tano issues bonds with a par value of $80,000 on January 1, 2017. The bonds’ annual contract rate is 8%, and interest is paid semiannually on June 30 and December 31. The bonds mature in three years. The annual market rate at the date of issuance is 10%, and the bonds are sold for $75,938. 1. What is the amount of the discount on these bonds at issuance? 2. How much total bond interest expense will be recognized over the life of these bonds? 3. Prepare an amortization table using the straight-line method to amortize the discount for these bonds.
Answer #1
1
Discount 4062 =80000-75938
2
Total interest expense over life of bonds
6 payments of $ 3200 19200
Par value at maturity 80000
Total repaid 99200
Less: Amount borrowed 75938
Total bond interest expense 23262
3
Semiannual Interest period end Unamortized Discount Carrying value
01/01/2017 4062 75938
06/30/2017 3385 76615
12/31/2017 2708 77292
06/30/2018 2031 77969
12/31/2018 1354 78646
06/30/2019 677 79323
12/31/2019 0 80000
Workings:
Semiannual Interest payment 3200 =80000*8%/2
Semiannual Discount amortization 677 =4062/6